# Robust Design and Evaluation of

### v5

@@ Appendix D Connections to alternative sensitivity analysis models
for all $y^{*},\tilde{y}^{*}\in\{0,1\}$ and with probability one. Notice that $\Gamma=1$ again nests unconfoundedess. [^14] We refer the reader to Section 7 of ZhaoSmallBhattacharya(19) for an in-depth comparison of the marginal sensitivity model and Rosenbaum’s $\Gamma$-sensitivity model., [^15] See also AronowLee(2013) and MiratrixWagerZubizaretta(18) which also use a version of Rosenbaum’s $\Gamma$-sensitivity model to construct bounds on a finite-population from a random sample with unknown selection probabilities. yadlowsky2018bounds applies Rosenbaum’s $\Gamma$-sensitivity analysis model to derive bounds on the conditional average treatment effect and the average treatment effect.
Proposition D.2.
Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Rosenbaum’s sensitivity analysis model (19) for some $\Gamma>1$. Then $P(\cdot)$ satisfies Equation (6) with $\underline{\delta}(x;\eta)=(\Gamma^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$.
@@
For brevity, we omit the conditioning on $X_{i}$ throughout the proof. As a first step, apply Bayes’ rule and observe that
$\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*})\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*})\mathbb{P}(D_{i}=1)}.$
Then, further notice that
$\frac{\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=1)}=\frac{\sum_{y\in\{0,1\}}P(D_{i}=0\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}{\sum_{y\in\{0,1\}}P(D_{i}=1\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}$
Letting $y^{*}=\arg\max_{y^{*}\in\{0,1\}}\frac{P(D_{i}=0\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}{P(D_{i}=1\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}$, the quasi-linearity of the ratio function implies that
@@
This then implies that, for any $y\in\{0,1\}$,
@@
where the last inequality is implied by Rosenbaum’s sensitivity analysis model (19). From this, we follow the same argument as the proof of Proposition D.1 to show that $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$. The proof for the lower bound follows an analogous argument. ∎
Our first result relates to Lemma 2.2 in yadlowsky2018bounds, which shows that a version of Rosenbaum’s sensitivity analysis model implies a bound on $\mathbb{E}[Y_{i}^{*}\mid D_{i}=0,X_{i}]$ via the solution to an estimating equation for general outcomes.
Partial $c$-dependence:
A recent strand of the econometrics literature proposes the partial c-dependence model for conducting sensitivity analyses on unobserved confounding in causal inference (e.g., MastenPoirier(18); MastenPoirier(20)). Partial $c$-dependence specifies that for some $c>0$,
@@
for each $x\in\mathcal{X}$. The next result relates the Equation (6) to the partial c-dependence model.
Proposition D.3.
$(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies, for all $x\in\mathcal{X}$,
$\left|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}-x)\right|\leq C$
if and only if
$\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c(x)$
for all $x\in\mathcal{X}$, where $c(x)=C\max\{\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)},\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=0\mid X_{i}=x)}\}$.
@@
To show this result, we first rewrite the bound
$|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)|\leq C$
as
$|\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1,X_{i}=x)\mathbb{P}(D_{i}=1)-\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i}=x)\mathbb{P}(D_{i}=0)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}.$
We can further rewrite the right hand side to arrive at
$|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}.$
Analogously, we can rewrite the bound as
$|\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=1,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=0,X_{i}=x)|\leq C,$
which in turn implies
$|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=0,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=0)}.$
by the same argument. ∎
Tukey’s factorization:
@@
Finally, a large literature conducts sensitivity analyses in missing data problems via Tukey’s factorization or outcome modeling. A popular approach specifies flexible parametric models for the difference between the unobserved conditional distribution $Y_{i}^{*}\mid\{X_{i},D_{i}=0\}$ and the observed conditional distribution $Y_{i}^{*}\mid\{X_{i},D_{i}=1\}$, or between $Y_{i}^{*}\mid\{X_{i},D_{i}=0\}$ and $Y_{i}^{*}\mid X_{i}$ (e.g., RotnitzkyEtAl(01); BirminghamEtAl(03); brumback2004sensitivity). For example, RobinsEtAl(00); FranksEtAl(19); Scharfstein(21) consider a sensitivity analysis model that assumes $\mathbb{P}(Y_{i}^{*}\mid D_{i}=0,X_{i})=\mathbb{P}(Y_{i}^{*}\mid D_{i}=1,X_{i})\frac{\exp(\gamma_{t}s_{t}(Y_{i}^{*})}{C(\gamma_{t};X_{i})}$, where $\gamma_{t}$ is a chosen parameter and $s_{t}(\cdot)$ is a chosen “tilting function.” For particular fixed choices of $\gamma_{t}$, $s_{t}(\cdot)$, such a model is sufficient to point identify various quantities of interest such as the conditional probability and the predictive performance estimands we consider. The literature then recommends that researchers report a sensitivity analysis that summarizes how their conclusions vary for alternative choices of $\gamma_{t}$ or $s_{t}(\cdot)$. In practice, however, it may be difficult, for the user to specify domain-specific knowledge that completely summarizes the relationship between these conditional distributions. Furthermore, any particular choice of the sensitivity analysis parameter $\gamma_{t}$ and tilting function $s_{t}(\cdot)$ may be mis-specified, and it is common that users only report results for a few choices. This outcome modeling approach may not encompass all possible values of the unidentified quantities that are consistent with the user’s domain-knowledge.
@@
An alternative approach places bounds on the mean difference in potential outcomes under treatment and control luedtke2015statistics; diaz2013sensitivity; diaz2018sensitivity. Our identification framework extends this approach by placing bounds on the covariate-conditional difference in means. We consider all joint distributions $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ that are consistent with the observable data and the researcher-specified bounds on the mean difference $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$.
Appendix E Additional Tables for Monte Carlo Simulations and Empirical Application
In this section, we collect together additional tables from the Monte Carlo simulations and empirical application that are referenced in the main text. Appendix Table 2b reports the average bias of our estimators for the overall performance bounds and the estimated coverage rate of 95% nominal confidence intervals based on Proposition 4.2 for $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$ and focusing on the lower bound on mean square error and the upper bound on accuracy. Table 3 provides detailed descriptions of the variable names in right panel of Figure 3b in Section 6.1 of the main text. Personal loans applications can have multiple listed applicants, so some variables refer to just the first listed applicant.

### v6

@@ E.2 Rosenbaum’s $\Gamma$-Sensitivity Model
for all $y^{*},\tilde{y}^{*}\in\{0,1\}$ and with probability one. Notice that $\Gamma=1$ again nests unconfoundedess.
AronowLee(2013) and MiratrixWagerZubizaretta(18) use a version of Rosenbaum’s $\Gamma$-sensitivity model to construct bounds on a finite-population from a random sample with unknown selection probabilities. yadlowsky2018bounds applies Rosenbaum’s $\Gamma$-sensitivity analysis model to derive bounds on the conditional average treatment effect and the average treatment effect. We refer the reader to Section 7 of ZhaoSmallBhattacharya(19) for an in-depth comparison of the marginal sensitivity model and Rosenbaum’s $\Gamma$-sensitivity model.
We can relate Rosenbaum’s $\Gamma$-sensitivity model to Assumption 2.1 under observed outcome bounds.
Proposition E.2.
Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Rosenbaum’s sensitivity analysis model (16) for some $\Gamma>1$. Then $P(\cdot)$ satisfies Assumption 2.1 with $\underline{\delta}(x;\eta)=(\Gamma^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$.
@@
For brevity, we omit the conditioning on $X_{i}$ throughout the proof. As a first step, we again apply Bayes’ rule and observe that $\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*})\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*})\mathbb{P}(D_{i}=1)}$. Then, further notice that $\frac{\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=1)}=\frac{\sum_{y\in\{0,1\}}P(D_{i}=0\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}{\sum_{y\in\{0,1\}}P(D_{i}=1\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}.$ Letting $y^{*}=\arg\max_{y^{*}\in\{0,1\}}\frac{P(D_{i}=0\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}{P(D_{i}=1\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}$, the quasi-linearity of the ratio function implies that
@@
This implies, for any $y\in\{0,1\}$,
@@
where the last inequality is implied by Rosenbaum’s sensitivity analysis model (16). From this, we follow the same argument as the proof of Proposition E.1 to show that $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$. The proof for the lower bound follows an analogous argument. ∎
Rosenbaum’s $\Gamma$-sensitivity model with parameter implies observed outcome bounds with $\underline{\Gamma}=\Gamma^{-1}$ and $\overline{\Gamma}=\Gamma$. However, the reverse direction does not hold — observed outcome bounds are weaker than Rosenbaum’s model.
E.3 Partial $c$-Dependence
A recent strand of the econometrics literature proposes the partial c-dependence model for conducting sensitivity analyses on unobserved confounding in causal inference (e.g., MastenPoirier(18); MastenPoirier(20)). Cast into our setting, partial $c$-dependence specifies that, for some $c>0$,
@@ E.3 Partial $c$-Dependence
for all $x\in\mathcal{X}$. Partial $c$-dependence bounds how much observing the outcome changes the selection problem. When $c=0$, we have unconfoundedness. The parameter $c$ represents the maximum percentage point difference in selection probabilities between different outcome values. The next result shows that Assumption 2.1 and partial c-dependence model can be related to one another through Bayes’ rule.
Proposition E.3.
  1. (i)
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies $\left|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)\right|\leq C$ for all $x\in\mathcal{X}$. Then, $\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c(x)$ for all $x\in\mathcal{X}$, where $c(x)=C\ \max\{\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)},\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=0\mid X_{i}=x)}\}$.
  2. (ii)
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies $\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c$ for all $x\in\mathcal{X}$. Then, $\left|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)\right|\leq c\ \min\{\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)},\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=0\mid X_{i}=x)}\}$ for all $x\in\mathcal{X}$.
@@
For brevity, we omit the conditioning on $X_{i}$ throughout the proof. To show the first result, we first rewrite $|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)|\leq C$ as $|\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)-\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}$. We can further rewrite the left hand side to arrive at $|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}$. Analogously, we can rewrite the bound as $|\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=1,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=0,X_{i}=x)|\leq C$, which in turn implies $|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=0,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=0)}$ by the same argument. The second result follows by the same argument. ∎
As mentioned, partial $c$-dependence has an interpretation in terms of percentage point changes in selection probabilities, while observed outcome bounds has an interpretation in terms of changes in outcome probabilities. In applications like pretrial release or consumer lending, the appropriate value of $c$ may be challenging to calibrate — it requires articulating assumptions about how judge and lender decision rates differ across defendants and applicants who would versus would not experience adverse outcomes. Observed outcome bounds ask researchers to directly articulate assumptions about outcome rates directly (e.g., ”rejected applicants are at most twice as likely to default”). This may align more naturally with available domain knowledge or proxy data.
Of course, researchers should choose the parameterization that best aligns with their substantive knowledge and available evidence. Moreover, Proposition E.3 provides explicit formulas that allow researchers to translate between the two frameworks, enabling those who find it easier to specify partial $c$-dependence to map to implied observed outcome bounds (and vice versa).
E.4 Tukey’s Factorization and Outcome Models
@@
A alternative approach to sensitivity analysis specifies outcome models via Tukey’s factorization that directly parameterize the relationship between $\mathbb{P}(Y^{*}_{i}\mid D_{i}=0,X_{i})$ and $\mathbb{P}(Y^{*}_{i}\mid D_{i}=1,X_{i})$ or $\mathbb{P}(Y^{*}_{i}\mid X_{i})$ (e.g., RotnitzkyEtAl(01); BirminghamEtAl(03); brumback2004sensitivity). For example, RobinsEtAl(00); FranksEtAl(19); Scharfstein(21) assume $\mathbb{P}(Y_{i}^{*}\mid D_{i}=0,X_{i})=\mathbb{P}(Y_{i}^{*}\mid D_{i}=1,X_{i})\frac{\exp(\gamma_{t}s_{t}(Y_{i}^{*}))}{C(\gamma_{t};X_{i})}$, where $\gamma_{t}$ is a chosen parameter and $s_{t}(\cdot)$ is a chosen “tilting function.” For particular fixed choices of $\gamma_{t}$, $s_{t}(\cdot)$, such a model is sufficient to point identify various quantities of interest such as the conditional probability and the predictive performance estimands we consider.
The key difference relative to our framework is how uncertainty is communicated. Researchers in this literature report how conclusions vary for alternative choices of $\gamma_{t}$ or $s_{t}(\cdot)$. By contrast, we suggest that researchers should articulate bounding functions that encode their beliefs about unobserved confounding, and then report the fully identified set consistent with those bounds. These bounding functions can formalize common heuristics without requiring a fully specified outcome model that can be difficult to justify.
@@
An alternative approach places bounds on the mean difference in potential outcomes under treatment and control luedtke2015statistics; diaz2013sensitivity; diaz2018sensitivity. This is analogous to the “approximate mean independence” assumption in Manski(03). Our identification framework extends this approach by placing bounds on the covariate-conditional difference in means.
Appendix F Monte Carlo Simulations
In this section, we illustrate that the finite sample behavior of our proposed estimators. We simulate training and evaluation datasets satisfying Assumption 2.1 under observed outcome bounds. We draw $X_{i}=\left(X_{i,1},\ldots,X_{i,d}\right)^{\prime}\sim N(0,I_{d})$ and $D_{i}$ conditional on $X_{i}$ according to $\mathbb{P}(D_{i}=1\mid X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\pi}}}\sum_{d=1}^{d_{\pi}}X_{i,d}\right)$ for some $d_{\pi}\in\{1,\ldots,d\}$ and $\sigma(a)=\frac{\exp(a)}{1+\exp(a)}$. We finally draw $Y_{i}^{*}$ conditional on $(D_{i},X_{i})$ according to
for some $d_{\mu}\in\{1,\ldots,d\}$ and $0<\Gamma_{true}<1$ We set $d=50$, $d_{\pi}=20$, $d_{\mu}=25$, and $\Gamma_{true}=0.75$.
F.1 Behavior of Estimated Conditional Probability Bounds
We compare three estimators for the upper bound on the conditional likelihood $\overline{\mu}^{*}(\cdot)$ with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$: first, our proposed estimators; second, the infeasible oracle that uses the true observed conditional probability and propensity score; and finally, a plug-in learner that does not use sample splitting nor pseudo-outcomes. Our estimator and the oracle procedure are constructed using a single split of the evaluation data. The first-stage nuisance functions $\eta=(\pi_{1}(\cdot),\mu_{1}(\cdot))$ are estimated using cross-validated Lasso logistic regressions, and the second-stage nonparametric regression uses cross-validated Lasso. Proposition 3.1 established that the integrated mean-square error of our estimators converge at fast rates to the integrated mean-square error of the infeasible oracle. By contrast, the plug-in estimator will inherit errors from the estimation of the nuisance functions.
Figure A1: Average integrated mean square of our estimator, the oracle learner, and the plug-in learner for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$.
[…]
F.2 Behavior of Estimated Predictive Performance Bounds
To evaluate our estimators for the bounds on positive class performance and overall performance, we simulate a training dataset $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n_{train}$ and estimate a predictive algorithm $s(\cdot)$ that predicts $Y_{i}=1$ only on the selected data $D_{i}=1$. We evaluate the constructed predictive algorithm’s positive class performance $\mbox{perf}_{+}(s;\beta)$ and overall performance $\mbox{perf}(s;\beta)$ using evaluation datatsets $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ simulated from the same data generating process and observed outcome bounds for alternative choices of $\underline{\Gamma},\overline{\Gamma}$.
$n$ Estimate SD Bias
500 0.545 0.016 0.001
1000 0.544 0.010 0.000
2500 0.544 0.007 0.000
Table A1: Bias properties for estimator for the upper bound on the true positive rate.
We explore how well our proposed estimator recover the bound on the true positive rate $\overline{\mbox{perf}}_{+}(s;\beta)$ across $1000$ simulated datasets of varying size $n\in\{500,1000,2500\}$ and a fixed choice $\underline{\Gamma}=2/3$ and $\overline{\Gamma}=3/2$. We construct our estimator using three splits of the evaluation data. We estimate the first-stage nuisance functions using random forests in the first fold, and the estimated bounding functions using cross-validated Lasso in the second fold. Online Appendix Table A1 reports the average bias, and Online Appendix Figure A2 visualizes the density of our estimates across simulations. Their bias quickly converges to zero as the size of the evaluation data grows large. We further explore how the performance of proposed estimators varies as we vary the researcher’s assumptions on the strength of unobserved confounding. Online Appendix Figure A3 reports the simulation distribution of our proposed estimators for both the upper and lower bounds as we vary $\overline{\Gamma}=\widetilde{\Gamma}$ for $\widetilde{\Gamma}\in\{1,\ldots,2.5\}$ and set $\underline{\Gamma}=1/\widetilde{\Gamma}$.
Figure A2: Distribution of estimator for the upper bound on the true positive rate across Monte Carlo simulations with observed outcome bounds
Figure A3: Estimated bounds on the true positive rate as $\underline{\Gamma},\overline{\Gamma}$ varies.
Our proposed estimators for the overall performance bounds follow a standard debiased construction. We briefly illustrate that our proposed estimators for the overall performance bounds also converge quickly and confidence intervals based on the asymptotic normal approximation have good coverage properties. For $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$ and focusing on the lower bound on mean square error and the upper bound on accuracy, Online Appendix Table A2 reports the average bias of our estimators and the estimated coverage rate of 95% nominal confidence intervals. Our proposed estimators are approximately unbiased for the true bounds. Their estimated standard errors slightly underestimate the true standard errors when the size of the evaluation dataset is small, but are quite accurate for $n\geq 1000$. Consequently, confidence intervals based on the asymptotic normal approximation have approximately 95% coverage. Our proposed estimators are approximately normally distributed in finite samples and concentrate around the true bounds (Online Appendix Figure A4).